OpenStudy (anonymous):

Consider the solid S described below. The base of S is a circular disk with radius 2r. Parallel cross-sections perpendicular to the base are squares. Find the volume V of this solid.

OpenStudy (dumbcow):

Equation of a circle with radius of 2r \[x^{2} + y^{2} = 4r^{2}\] solving for y \[y = \sqrt{4r^{2} - x^{2}}\] Each square has side length of 2y Area of each square is (2y)^2 = 4y^2 Its best to integrate with respect to x, with bounds from -2r to 2r so change area so its in terms of x \[4y^{2} = 16r^{2} - 4x^{2}\] \[V = \int\limits_{-2r}^{2r} 16r^{2} -4x^{2} dx\]